# ChatModel: Automating Reference Model Design and Verification with LLMs

### v2

@@ 1 Introduction
The rapidly increasing complexity of integrated circuit (IC) designs has significantly prolonged functional verification cycles, posing a critical bottleneck in agile hardware development [37, 21, 32]. In this context, model-driven verification (MDV) is recognized as a highly effective methodology and has been widely applied in preliminary chip designs, such as neural processors and artificial intelligence accelerators [66, 14, 52, 43, 22, 33]. The models in MDV are typically written in high-level abstraction languages (i.e. Python [29, 3, 48], C++ [70, 67], MATLAB [19, 74], or SystemC [7, 1, 50, 2]) by engineers and serve as golden references for hardware designs. As shown in Figure 2, by systematically comparing the model outputs with those of the Device Under Test (DUT) under identical input stimuli, developers can effectively identify design flaws and accelerate the debugging and iteration process [53].
@@
In the MDV framework, open-source simulators or loosely-timed modeling approaches are often employed to rapidly develop models [6, 23, 8, 5]. However, the execution processes within these models exhibit substantial discrepancies compared to actual hardware execution, thereby restricting their effectiveness in diagnosing errors within the DUT. For the development of specialized designs, developers must consider timing factors to construct more accurate models. However, building high-fidelity reference models requires a deep understanding of the behavior and verification requirements of the target hardware. As hardware designs grow increasingly intricate, the expertise and effort required expand exponentially, hindering agile model development.
@@
With the advent of large language models (LLMs), a revolutionary paradigm shift has occurred in the field of natural language processing (NLP) [47]. Recent studies have showcased their impressive capabilities in code generation for programming languages [85, 56, 28, 44, 77, 90] such as C++ and Python, as well as hardware description languages [79, 16, 78, 75] such as Verilog. However, their application in constructing reference models remains underexplored. General-purpose LLMs encounter numerous challenges when applied to reference modeling (detailed further in Appendix A), including context and token limitations, hallucinations, and limited self-correction capabilities. These issues undermine the quality and stability of the outputs, restricting LLMs to small-scale designs. Moreover, the increasing diversity of design specifications and hardware architectures further exacerbates these issues, making it challenging for LLMs to meet the demands of constructing reference models.
@@
To overcome these limitations, we introduce ChatModel, an end-to-end platform designed for agile generation and verification of reference models, as illustrated in Figure 2. The ChatModel platform leverages a multi-agent LLM system comprising two integrated groups: one standardizes design specifications, while the other automates the generation and validation of reference models. By adaptively decomposing complex design tasks into specialized subtasks, ChatModel facilitates the rapid construction of corresponding design Intermediate Representations (IR) and their associated reference models. Consequently, the ChatModel enables agile reference modeling, thereby substantially reducing both the time and effort required for the model design and validation. The contributions of our work are summarized as follows: In the first phase, specification standardization agents group automates the standardization process based on design specifications, transforming non-standard specifications into a standardized Intermediate Representation (IR). This IR functions as a novel domain-specific language for describing reference models. In the second phase, reference model generation agents group utilizes the generated Design IR to implement the Hierarchical Agile Modeling (HAM) flow for reference model generation. HAM employs an adaptive task planning algorithm to devise the construction strategy for the reference model, facilitating automated generation and verification from module to subsystem to system in a “building block” manner.
@@ 2 Related Works
Traditional Reference Model Design: As a vital component in the verification process, reference models are typically developed using programming languages at varying abstraction levels, tailored to the functionality of the DUT [65, 54]. For image or signal processing designs, Python [68, 62, 4] and MATLAB [38, 91] are commonly employed for modeling, while C++ [26, 51, 67] is predominantly used for constructing reference models for logic functions and CPU designs. Recently, SystemC with Transaction-Level Modeling 2.0 (TLM 2.0) has achieved widespread adoption for reference model design, attributable to its capacity to facilitate modeling across multiple levels of abstraction [25, 31, 59, 39, 17]. For larger System-on-Chip (SoC) designs, simulators like gem5 [9, 63, 72] or SST [69] are employed for model development to verify hardware system correctness, but their limited configurability results in coarse simulation granularity. Although these approaches have contributed to the advancement of reference model development, they require developers to possess substantial expertise in hardware and software programming, often necessitating a considerable and sustained investment of resources and time.
@@
LLM in Code Generation: With the emergence of models such as ChatGPT [35] and Claude [10], LLMs have demonstrated their powerful capabilities in automating code generation from natural language descriptions [15, 88, 92, 41]. Recent efforts [73, 55, 12, 78, 11, 18, 82] have explored LLM-aided Hardware Description Language (HDL) generation and debugging using prompt engineering [20, 86, 57, 76], retrieval-augmented generation [46, 61], and supervised fine-tuning [36, 87, 83]. Besides HDL generation, some works have also investigated complex software designs. The learning selection method LAIL [40] and the self-planning code generation technique [30] are both enhancements in these areas. A two-stage code generation method [89] has demonstrated improved efficiency in generating code under zero-shot conditions. The LCG model [42], which adopts a multi-LLM agent cooperation mechanism, integrates seamlessly into traditional software development processes, significantly elevating the quality of generated code. In addition, study like UVLLM [27] has explored the potential of LLM in the generation of reference models. However, implementations rely on the relatively lower-level language SystemVerilog, which restricts the scale of the generated models. Despite these advancements, the potential of LLMs to generate reference models for large-scale hardware designs remains underexplored, with notable limitations in both the token/context memory limits for complex designs and limited debugging capabilities of LLMs.
@@ 5.1 Design Planning
1 Function Build_Seq($D_{\text{tree}}$, node, visited):
@@ 5.3 Model Verification & Debugger
To enhance the debugging capabilities of LLMs, we document common errors and their solutions within the reference model and utilize Chroma [64] to construct a vector database. Agent 7 employs Retrieval-Augmented Generation (RAG) technology to access relevant data from this database, facilitating the resolution of syntactic and functional errors during validation. For syntactic errors, terminal output logs are used to retrieve pertinent information from documents or knowledge bases via RAG technology. This information, along with relevant code snippets, is input into the LLM for error analysis and code correction. In the case of functional errors, Agent 7 also extracts logs from the period surrounding the error occurrence to enable the LLM to analyze the causes of output discrepancies.
@@ 6 Evaluations
Zero-Shot [84] 68.21 42.15 56.73 22.39 46.91 19.36 59.62 31.72
Few-Shot [80] 75.34 48.07 66.31 24.24 49.23 22.32 65.65 36.03
CoT [81] 73.47 51.44 67.14 27.37 53.82 23.11 66.26 38.55
CoT+Few-Shot [34] 87.41 53.21 71.04 31.53 55.14 28.31 71.45 41.79
Multi-Agent [58] 87.87 61.56 73.17 34.54 58.38 31.14 76.18 47.53
Best-Choice [60] 97.43 57.28 70.82 33.31 54.61 28.37 79.58 44.23
@@ 6.1 Experimental setup
Benchmarks. To evaluate the effectiveness of our work, we employed 300 designs from the open-source benchmarks VerilogEval [45] and RTLLM2.0 [49], as well as our own ModelEval dataset. Both VerilogEval and RTLLM2.0 are relatively simple and primarily used to evaluate the generative capabilities of LLMs, yet their design specifications can still be applied to generating reference models. To more accurately evaluate the capability of ChatModel in generating complex reference models, we developed the ModelEval benchmark, which consists of 94 distinct designs that vary significantly in complexity. Each design includes comprehensive test cases to verify the functional correctness of the generated reference models. The model is considered successfully generated only if it passes these test cases.
Evaluation Metrics. Since LLMs are probabilistic, each case is evaluated multiple times to account for output variability, and the pass@k [13] metric is calculated. Unless otherwise specified, ChatModel uses GPT-4o as the foundational LLM for each agent, and the temperature of all LLMs is set to 0.3 to balance generation capability and randomness. The debugging iteration threshold in ChatModel is set at 6, as improvements beyond this point are difficult to observe. For more detailed information about the experiment setup, please see Appendix B.
@@ 6.2 Capability for Reference Model Generation
Additionally, we performed five repeated tests on all benchmarks with various k values, as shown in Figure 6 (b) and (c). The results indicate that ChatModel not only enhances the ability of LLMs to generate reference models, but also significantly improves the stability and robustness of the outputs. Under the pass@1 condition, ChatModel achieves a 51.62% improvement in the FP rate compared to the Zero-Shot baseline, with variability controlled at 7.45%. This suggests that the majority of tests consistently pass on the first attempt. Moreover, as the value of k increases, the improvement in the false positive rate increases to 55.02%, while the variability decreases to 4.08%.
@@ Appendix A Challenges in Reference Model Generation with LLMs
Pass@$k$ $:=\mathop{\mathbb{E}}_{\text{Problems}}\left[1-\frac{{\binom{n-c}{k}}}{\binom{n}{k}}\right]$ (1)
@@
Here, Pass@k [13] is a widely used metric for assessing the generative capabilities of LLMs. This metric involves generating n code samples for each task (where n is greater than k), and then randomly selecting k samples. The test is deemed successful if at least one of these k samples passes the unit tests. The variable c represents the total number of samples that successfully pass the unit tests. In the formula of $D_{C}$, N denotes the number of debugging iterations required during functional testing, while $Err_{i}$ indicates the number of errors encountered in each debugging iteration. A higher score according to this formula signifies superior performance of the LLM.
@@ B.1 Benchmarks
In this section, we provide a comprehensive overview of the benchmarks utilized in our evaluation. Owing to the limited availability of open-source benchmarks specifically designed for reference models of hardware design, we have adopted VerilogEval and RTLLM2.0 for our experiments. Although these benchmarks primarily aim to assess LLMs in generating hardware description languages (HDLs), the provided design specifications can also be applied to generate SystemC reference models. By utilizing EDA tools such as VCS [71], the generated models can be co-simulated with the supplied test platforms to verify their functional equivalence to the hardware designs. Furthermore, we developed a benchmark called ModelEval to more accurately evaluate LLM performance under complex design requirements.
VerilogEval [45] is a benchmark developed by NVIDIA to evaluate the performance of LLMs in generating Verilog code for hardware design and validation. It includes 156 diverse generation tasks, spanning from simple logic circuits to complex finite state machines. As shown in the Table 2, each design is provided with four files: interface, specification, reference design, and testbench.
@@
RTLLM2.0 [49] introduced by the Hong Kong University of Science and Technology, serves as a benchmark to assess the ability of LLMs to generate HDLs from natural language. This benchmark includes 50 diverse functional designs, categorized into four types: arithmetic, memory, control, and miscellaneous, with each design containing all the information specified in Table 2.
@@ Appendix C Supplementary Experiment
To further evaluate the impact of various large language models on ChatModel and other LLM-based methodologies, we selected Claude 3.5-Sonnet [10] and DeepSeek-R1 [24] for testing. We employed the single-agent approach (CoT+Few-Shot) and the multi-agent approach (Multi-Agent) from Table 1, which exhibited superior performance as benchmarks for comparison with ChatModel. The experimental results, presented in Table 5, indicate that ChatModel substantially improves functionality pass rates, with a 48.81% increase when using Claude 3.5-Sonnet and a 37.01% improvement with DeepSeek-R1, relative to the Zero-Shot baseline. These findings suggest that the framework consistently exhibits robust performance and significant enhancement when integrated with different LLMs.

### v3

@@ 1 Introduction
The rapidly increasing complexity of integrated circuit (IC) designs has significantly prolonged functional verification cycles, posing a critical bottleneck in agile hardware development lahti2018we ; bottleneck2 ; logicfuzzer . In this context, model-driven verification (MDV) is recognized as a highly effective methodology and has been widely applied in preliminary chip designs, such as neural processors and artificial intelligence accelerators quinones2020ampere ; cirstea2024digital ; mishty2024ai ; linehan2012aspect ; genius2017model ; kaja2024advanced . The models in MDV are typically written in high-level abstraction languages (i.e. Python jiang2020pymtl3 ; batten2018open ; logaras2014python , C++ shcherbakov2011bringing ; raia2024case , MATLAB delva2008using ; versen2016model , or SystemC bergeron2012writing ; atac2014hdl ; mathur2007design ; balasubramanian2023bit ) by engineers and serve as golden references for hardware designs. As shown in Figure 2, by systematically comparing the model outputs with those of the Device Under Test (DUT) under identical input stimuli, developers can effectively identify design flaws and accelerate the debugging and iteration process mohamed2016ip .
@@
In the MDV framework, open-source simulators or loosely-timed modeling approaches are often employed to rapidly develop models bellard2005qemu ; ghosh2015case ; bidmeshki2017information ; becker2015challenges . However, the execution processes within these models exhibit substantial discrepancies compared to actual hardware execution, thereby restricting their effectiveness in diagnosing errors within the DUT. For the development of specialized designs, developers must consider timing factors to construct more accurate models. However, building high-fidelity reference models requires a deep understanding of the behavior and verification requirements of the target hardware. As hardware designs grow increasingly intricate, the expertise and effort required expand exponentially, hindering agile model development.
@@
With the advent of large language models (LLMs), a revolutionary paradigm shift has occurred in the field of natural language processing (NLP) liu2023evaluating . Recent studies have showcased their impressive capabilities in code generation for programming languages xu2022systematic ; ni2024l2ceval ; jiang2024survey ; liu2023your ; wang2023review ; zheng2024towards such as C++ and Python, as well as hardware description languages wang2024chatcpu ; danneels2024exploring ; wang2024large ; wan2025genben such as Verilog. However, their application in constructing reference models remains underexplored. General-purpose LLMs encounter numerous challenges when applied to reference modeling (detailed further in Appendix A), including context and token limitations, hallucinations, and limited self-correction capabilities. These issues undermine the quality and stability of the outputs, restricting LLMs to small-scale designs. Moreover, the increasing diversity of design specifications and hardware architectures further exacerbates these issues, making it challenging for LLMs to meet the demands of constructing reference models.
@@
To overcome these limitations, we introduce ChatModel, an end-to-end platform designed for agile generation and verification of reference models, as illustrated in Figure 2. The ChatModel leverages a multi-agent LLM system comprising two integrated groups: one standardizes design specifications, while the other automates the generation and validation of reference models. By adaptively decomposing complex design tasks into specialized subtasks, ChatModel facilitates the rapid construction of the corresponding design Intermediate Representations (IR). The Design IR serves as the input to the Hierarchical Agile Modeling (HAM) flow to generate the reference model. By employing an adaptive task planning algorithm, HAM facilitates automated and scalable generation and verification of reference models, seamlessly progressing from module to subsystem to system in a “building block” manner.
@@ 2 Related Works
Traditional Reference Model Design: As a vital component in the verification process, reference models are typically developed using programming languages at varying abstraction levels, tailored to the functionality of the DUT qamar2020comprehensive ; moursi2018different . For image or signal processing designs, Python renuka2023monitoring ; platonova2024using ; beaulieu2017co and MATLAB lakshminarayana2022flattening ; zheng2024generic are commonly employed for modeling, while C++ heikura2021high ; mintz2006hardware ; raia2024case is predominantly used for constructing reference models for logic functions and CPU designs. Recently, SystemC with Transaction-Level Modeling 2.0 (TLM 2.0) has achieved widespread adoption for reference model design, attributable to its capacity to facilitate modeling across multiple levels of abstraction habibi2006design ; jindal2003verification ; park2008co ; le2016towards ; delbergue2016qbox . For larger System-on-Chip (SoC) designs, simulators like gem5 binkert2011gem5 ; power2014gem5 ; ta2018simulating or SST rodrigues2011structural are employed for model development to verify hardware system correctness, but their limited configurability results in coarse simulation granularity. Although these approaches have contributed to the advancement of reference model development, they require developers to possess substantial expertise in hardware and software programming, often necessitating a considerable and sustained investment of resources and time.
@@
LLM in Code Generation: With the emergence of models such as ChatGPT kocon2023chatgpt and Claude caruccio2024claude , LLMs have demonstrated their powerful capabilities in automating code generation from natural language descriptions coello2024effectiveness ; zhang2023planning ; zhong2024can ; li2025idse . Recent efforts thakur2024verigen ; nadimi2024multi ; chang2023chipgpt ; wang2024large ; chang2024data ; delorenzo2024make ; wong2024vgv have explored LLM-aided Hardware Description Language (HDL) generation and debugging using prompt engineering d2024exploring ; xu2025novel ; niu2025rechisel ; wan2024assessment , retrieval-augmented generation liu2024chatchisel ; ping2025hdlcore , and supervised fine-tuning kumar2024hdl ; yao2024hdldebugger ; wu2024itertl . Besides HDL generation, some works have also investigated complex software designs. The learning selection method LAIL li2023large and the self-planning code generation technique jiang2023self are both enhancements in these areas. A two-stage code generation method zhao2024two has demonstrated improved efficiency in generating code under zero-shot conditions. The LCG model lin2024llm , which adopts a multi-LLM agent cooperation mechanism, integrates seamlessly into traditional software development processes, significantly elevating the quality of generated code. In addition, study like UVLLM hu2024uvllm has explored the potential of LLM in the generation of reference models. However, implementations rely on the relatively lower-level language SystemVerilog, which restricts the scale of the generated models. Despite these advancements, the potential of LLMs to generate reference models for large-scale hardware designs remains underexplored, with notable limitations in both the token/context memory limits for complex designs and limited debugging capabilities of LLMs.
@@ 5.1 Design Planning
[…]
@@ 5.3 Model Verification & Debugger
To enhance the debugging capabilities of LLMs, we document common errors and their solutions within the reference model and utilize Chroma prabhune2024deploying to construct a vector database. Agent 7 employs Retrieval-Augmented Generation (RAG) technology to access relevant data from this database, facilitating the resolution of syntactic and functional errors during validation. For syntactic errors, terminal output logs are used to retrieve pertinent information from documents or knowledge bases via RAG technology. This information, along with relevant code snippets, is input into the LLM for error analysis and code correction. In the case of functional errors, Agent 7 also extracts logs from the period surrounding the error occurrence to enable the LLM to analyze the causes of output discrepancies.
@@ 6 Evaluations
Zero-Shot xian2018zero 68.21 42.15 56.73 22.39 46.91 19.36 59.62 31.72
Few-Shot wang2020generalizing 75.34 48.07 66.31 24.24 49.23 22.32 65.65 36.03
CoT wei2022chain 73.47 51.44 67.14 27.37 53.82 23.11 66.26 38.55
CoT+Few-Shot kim2023cot 87.41 53.21 71.04 31.53 55.14 28.31 71.45 41.79
Multi-Agent panait2005cooperative 87.87 61.56 73.17 34.54 58.38 31.14 76.18 47.53
Best-Choice partalas2007multi 97.43 57.28 70.82 33.31 54.61 28.37 79.58 44.23
@@ 6.1 Experimental setup
Benchmarks. To evaluate the effectiveness of our work, we employed 300 designs from the open-source benchmarks VerilogEval liu2023verilogeval and RTLLM2.0 lu2024rtllm , as well as our own ModelEval dataset. Both VerilogEval and RTLLM2.0 are relatively simple and primarily used to evaluate the generative capabilities of LLMs, yet their design specifications can still be applied to generating reference models. To more accurately evaluate the capability of ChatModel in generating complex reference models, we developed the ModelEval benchmark, which consists of 94 distinct designs that vary significantly in complexity. Each design includes comprehensive test cases to verify the functional correctness of the generated reference models. The model is considered successfully generated only if it passes these test cases.
Evaluation Metrics. Since LLMs are probabilistic, each case is evaluated multiple times to account for output variability, and the pass@k chen2021evaluating metric is calculated. Unless otherwise specified, ChatModel uses GPT-4o as the foundational LLM for each agent, and the temperature of all LLMs is set to 0.3 to balance generation capability and randomness. The debugging iteration threshold in ChatModel is set at 6, as improvements beyond this point are difficult to observe. For more detailed information about the experiment setup, please see Appendix B.
@@ 6.2 Capability for Reference Model Generation
Additionally, we performed five repeated tests on all benchmarks with various k values, as shown in Figure 6 (b) and (c). The results indicate that ChatModel not only enhances the ability of LLMs to generate reference models, but also significantly improves the stability and robustness of the outputs. Under the pass@1 condition, ChatModel achieves a 51.62% improvement in the FP rate compared to the Zero-Shot baseline, with variability controlled at 7.45%. This suggests that the majority of tests consistently pass on the first attempt. Moreover, as the value of k increases, the improvement in the FP rate increases to 55.02%, while the variability decreases to 4.08%.
@@ Appendix A Challenges in Reference Model Generation with LLMs
[…]
@@
Here, Pass@k chen2021evaluating is a widely used metric for assessing the generative capabilities of LLMs. This metric involves generating n code samples for each task (where n is greater than k), and then randomly selecting k samples. The test is deemed successful if at least one of these k samples passes the unit tests. The variable c represents the total number of samples that successfully pass the unit tests. In the formula of $D_{C}$, N denotes the number of debugging iterations required during functional testing, while $Err_{i}$ indicates the number of errors encountered in each debugging iteration. A higher score according to this formula signifies superior performance of the LLM.
@@ B.1 Benchmarks
In this section, we provide a comprehensive overview of the benchmarks utilized in our evaluation. Owing to the limited availability of open-source benchmarks specifically designed for reference models of hardware design, we have adopted VerilogEval and RTLLM2.0 for our experiments. Although these benchmarks primarily aim to assess LLMs in generating hardware description languages (HDLs), the provided design specifications can also be applied to generate SystemC reference models. By utilizing EDA tools such as VCS sutherland2006modeling , the generated models can be co-simulated with the supplied test platforms to verify their functional equivalence to the hardware designs. Furthermore, we developed a benchmark called ModelEval to more accurately evaluate LLM performance under complex design requirements.
VerilogEval liu2023verilogeval is a benchmark developed by NVIDIA to evaluate the performance of LLMs in generating Verilog code for hardware design and validation. It includes 156 diverse generation tasks, spanning from simple logic circuits to complex finite state machines. As shown in the Table 2, each design is provided with four files: interface, specification, reference design, and testbench.
@@
RTLLM2.0 lu2024rtllm introduced by the Hong Kong University of Science and Technology, serves as a benchmark to assess the ability of LLMs to generate HDLs from natural language. This benchmark includes 50 diverse functional designs, categorized into four types: arithmetic, memory, control, and miscellaneous, with each design containing all the information specified in Table 2.
@@ Appendix C Supplementary Experiment
To further evaluate the impact of various large language models on ChatModel and other LLM-based methodologies, we selected Claude 3.5-Sonnet caruccio2024claude and DeepSeek-R1 guo2025deepseek for testing. We employed the single-agent approach (CoT+Few-Shot) and the multi-agent approach (Multi-Agent) from Table 1, which exhibited superior performance as benchmarks for comparison with ChatModel. The experimental results, presented in Table 5, indicate that ChatModel substantially improves functionality pass rates, with a 48.81% increase when using Claude 3.5-Sonnet and a 37.01% improvement with DeepSeek-R1, relative to the Zero-Shot baseline. These findings suggest that the framework consistently exhibits robust performance and significant enhancement when integrated with different LLMs.